Circular No. 5799

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1993O AND 1993P IN ANONYMOUS GALAXIES M. Hamuy, Cerro Tololo Interamerican Observatory; and J. Maza,

University of Chile, report the discovery of an apparent supernova of m_{pg} ∼ 17.5 by R. Antezana (University of Chile), found on a 20-min unfiltered IIa-O plate taken on May 18.16 UT by G. Valladares with the CTIO Curtis Schmidt telescope. SN 1993O is located 14′′ west and 8′′ north of the nucleus of an anonymous galaxy at α = 13^{h}28^{m}19^{s}, δ = −32^{o}57′.6 (equinox 1950.0). Confirmation of this supernova was made by A. Layden (CTIO) on CCD B and V images obtained on May 20.229 with the CTIO 0.9-m telescope. M. M. Phillips (CTIO) obtained a spectrogram on May 21.13 with the CTIO 1.5-m telescope, indicating that SN 1993O is a type-Ia event at maximum light, or perhaps even a few days before maximum; a spectrum of the host galaxy shows weak Hα emission at a redshift of 0.051.

Phillips, Hamuy, and Maza also report the discovery of another supernova of m_{pg} ∼ 17.5 by Antezana on the same plate noted above, SN 1993P being located in the nucleus of a spiral galaxy at α = 13^{h}26^{m}37^{s}.4, δ = −30^{o}09′18′′ (equinox 1950.0). Confirmation was again made by Layden from a CCD V image obtained on May 20.22. A spectrogram acquired by Phillips on May 21.03 with the CTIO 1.5-m telescope reveals SN 1993P to be a type-Ic event near maximum light. Weak Hα emission observed in the disk of the galaxy yields a redshift of 0.048.

SUPERNOVA 1993J IN NGC 3031 D. A. Green and G. G. Pooley, Mullard Radio Astronomy Observatory,

Cambridge, communicate: “Continued monitoring of SN 1993J with the Ryle Telescope at 15.25 GHz (not 15 or 15.5 GHz, as we incorrectly stated on IAUC 5751 and 5773) shows that its increase in flux density at this frequency has recently slowed considerably. We see an approximately linear rise to 53 mJy on May 3.7 UT from its detection Apr. 5 (IAUC 5751; ∼ 10 days after the supernova event, taken to be Mar. 26.0), with a slower rise since then to 60 mJy by May 19.8. These flux densities are preliminary, and are based on an assumed flux density of 1.2 Jy for 0954+658. The linear increase in flux density seen at this frequency lasted ∼ 30 days, which is about three times the delay in detection of the supernova; this is difficult to reconcile with the available models for radio supernovae.”

1993 May 21 Daniel W. E. Green